OpenStudy (anonymous):

HELP IM RUNNING OUT OF TIME! The diameter of a cylinder is 8 cm and its height is 14 cm. Calculate the volume of the cylinder to the nearest cubic centimetre. Use π = 3.14.

OpenStudy (anonymous):

HELP!?

OpenStudy (capnkookcx):

I don't know :/

OpenStudy (anonymous):

AHHHHHHHH!!!!

OpenStudy (anonymous):

@Mehek14

OpenStudy (johnweldon1993):

\[\large V = \pi r^2 h\] \[\large V = 3.14 \times (4)^2 \times 14\] \[\large V = ?\]

OpenStudy (anonymous):

IDK!!!!

OpenStudy (anonymous):

174.84? @johnweldon1993

OpenStudy (johnweldon1993):

\[\large V = 3.14\times 16 \times 14\] \[\large V = 703.36\]